OpenStudy (anonymous):

The triangles shown are similar. What is the value of x? A. 12 B. 15 C. 10 D. 9

Mehek (mehek14):

\(\dfrac{8}{4}=\dfrac{x}{5}\)

Mehek (mehek14):

cross multiply

OpenStudy (anonymous):

so 8*5=40

pooja195 (pooja195):

good 4x=40 divide both sides by 4 to isolate x

pooja195 (pooja195):

4/4=? 40/4=?

OpenStudy (anonymous):

1/10

OpenStudy (anonymous):

so the answer is 10 or.....................???????????

pooja195 (pooja195):

4/4=x x=10

OpenStudy (anonymous):

so 10

pooja195 (pooja195):

so you had it somewhat correct 4/4=1 which would leave us with jut x you had the 10 correct